OpenStudy (anonymous):

determine the centroid of the surface of a right circular cone of altitude h

OpenStudy (anonymous):

The answer from the book is 2h/3 from the apex, but i dont know how it was solved

OpenStudy (anonymous):

Help please

OpenStudy (irishboy123):

are you using calculus? if so take a strip thickness dx and roll it out and you get a trapezoid as shown in blue  then calculate dM and dM_o for mass and moment and \(\large \bar x = \frac{\int\limits_{0}^{h} dM_o}{\int\limits_{0}^{h} dM}\) if you're not doing calculus, sorry :(